OpenStudy (gabylovesyou):

Given the function f(x) = 2(x + 6), find x if f(x) = 22. 56 5 38 17

OpenStudy (ghazi):

22=2(x+6) (x+6)=22/2=11...solve for x

OpenStudy (gabylovesyou):

5???

OpenStudy (ghazi):

yes